# Measuring and Controlling Solution Degeneracy across Task-Trained Recurrent Neural Networks

### v2

@@ 5 Acknowledgments
This work was supported by the NIH (RF1DA056403), James S. McDonnell Foundation (220020466), Simons Foundation (Pilot Extension-00003332-02), McKnight Endowment Fund, CIFAR Azrieli Global Scholar Program, NSF (2046583), and Harvard Medical School Dean’s Innovation Award. A.H. is supported by the Kempner Graduate Fellowship. We thank Blake Bordelon for insightful discussions on Maximal Update Parameterization, and we are grateful to Jiayi Zhang, Naomi Saphra, Arna Ghosh, Andy Keller, Yu Duan, Hanson Lu, and all members of the Rajan Lab for helpful discussions.
@@ B.4 Path Integration
Appendix C Memory demand of each task
@@
and train a two-layer MLP to predict the subsequent target $y_{t+1}$. We then evaluate the held-out mean-squared error $\mathrm{MSE}(h)$, averaged over multiple random initializations. We identify the smallest history length $h^{*}$ at which the error curve plateaus or has a minimum, and take $h^{*}$ as the task’s intrinsic memory demand.
@@
Figure 9: Memory demand of each task. The held-out mean-squared error $\mathrm{MSE}(h)$ of a two-layer MLP predictor is plotted against history length $h$. The intrinsic memory demand $h^{*}$, defined by the plateau or minimum of each curve, is 1 for the N-Bits Flip-Flop and Path Integration tasks, 2 for Sine Wave Generation, and 25 for Delayed Discrimination—matching the inter-stimulus delay interval in that task.
Appendix D More details on the degeneracy metrics
D.1 Dynamical Degeneracy
Briefly, DSA proceeds as follows: Given two RNNs with hidden states $\mathbf{h}_{1}(t)\in\mathbb{R}^{n}$ and $\mathbf{h}_{2}(t)\in\mathbb{R}^{n}$, we first generate a delay-embedded matrix, $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ of the hidden states in their original state space. Next, for each delay-embedded matrix, we use Dynamic Mode Decomposition (DMD) [73] to extract linear forward operators $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ of the two systems’ dynamics. Finally, a Procrustes distance between the two matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ is used to quantify the dissimilarity between the two dynamical systems and provide an overall DSA score, defined as:
@@ D.1 Dynamical Degeneracy
D.2 Weight degeneracy
We computed the pairwise distance between the recurrent matrices from different networks using Two-sided Permutation with One Transformation [74, 17] function from the Procrustes Python package [59].
Appendix E Representational degeneracy
@@
[…]
@@
Appendix F Detailed characterization of OOD generalization performance
@@
F.1 Changing task complexity
@@
F.2 Changing feature learning strength
@@
F.3 Changing network size
@@
F.4 Changing regularization strength
F.4.1 Low-rank regularization
@@ F.4.2 Sparsity (L1) regularization
Appendix G A short introduction to Maximal Update Parameterization ($\mu P$)
Under the NTK parametrization, as the network width goes to infinity, the network operates in the lazy regime, where its functional evolution is well-approximated by a first-order Taylor expansion around the initial parameters [39, 47, 13, 82]. In this limit feature learning is suppressed and training dynamics are governed by the fixed Neural Tangent Kernel (NTK).
To preserve non-trivial feature learning at large width, the Maximal Update Parametrization ($\mu$P) rescales both the weight initialisation and the learning rate. $\mu P$ keeps three quantities width-invariant at every layer—(i) the norm/variance of activations (ii) the norm/variance of the gradients, and (iii) the parameter updates applied by the optimizer [83, 84, 28, 7].
@@ Appendix G A short introduction to Maximal Update Parameterization ($\mu P$)
Appendix H Theoretical relationship between parameterizations
We compare two RNN formalisms used in different parts of the main manuscript: a standard discrete-time RNN trained with fixed learning rate and conventional initialization, and a $\mu$P-style RNN trained with leaky integrator dynamics and width-aware scaling.
@@ Appendix H Theoretical relationship between parameterizations
  • $\mu$P RNN: $\phi$ is applied pre-activation; i.e. before the recurrent weight matrix, so the hidden state is first non-linearized and then linearly combined
Miller and Fumarola [60] demonstrated that two classes of continuous-time firing-rate models which differ in their placement of the non-linearity are mathematically equivalent under a change of variables:
@@
This mapping applies directly to RNNs viewed as continuous-time dynamical systems and helps relate v-type $\mu$P-style RNNs to standard discrete-time RNNs. It suggests that the $\mu$P RNN (in v-type form) and the standard RNN (in r-type form) can be treated as different parameterizations of the same underlying dynamical system when:
@@
In summary, while a theoretical equivalence exists, it is contingent on consistent scaling across all components of the model. In this manuscript, we use the standard discrete-time RNNs due to its practical relevance for task-driven modeling community, while switching to $\mu P$ to isolate the effect of feature learning and network size. Additionally, we confirm that the feature learning and network size effects on degeneracy hold qualitatively the same in standard discrete-time RNNs, unless where altering network width induces unstable and lazier learning in larger networks (Figure L and M).
Appendix I Verifying larger $\gamma$ reliably induces stronger feature learning in $\mu P$
In $\mu P$ parameterization, the parameter $\gamma$ interpolates between lazy training and rich, feature-learning dynamics, without itself being the absolute magnitude of feature learning. Here, we assess feature-learning strength in RNNs under varying $\gamma$ using two complementary metrics:
@@ I.2 Delayed Discrimination
I.3 Sine Wave Generation
@@ I.4 Path Integration
Appendix J Verifying $\mu P$ reliably controls for feature learning across network width
Here, we only use Kernel Alignment to assess the feature learning strength in the networks since the unnormalized weight-change norm $\left\|\mathbf{W}_{T}-\mathbf{W}_{0}\right\|_{F}$ scales directly with matrix size (therefore network size) and there exists no obvious way to normalize across different dimensions. In our earlier analysis where we compared weight-change norms at varying $\gamma$, network size remained fixed, so those Frobenius-norm measures were directly comparable. We found that, for all tasks except Delayed Discrimination, the change in mean KA across different network sizes remains extremely small (less than 0.1), which demonstrates that $\mu P$ parameterization with the same $\gamma$ has effectively controlled for feature learning strength across network sizes. On Delayed Discrimination, the networks undergo slightly lazier learning for larger network sizes. Nevertheless, we still include Delayed Discrimination in our analyses of solution degeneracy to ensure our conclusions remain robust even when $\mu P$ can’t perfectly equalize feature-learning strength across widths. As shown in the main paper, lazier learning regime generally increases dynamical degeneracy; yet, larger networks which exhibit lazier learning in the N-BFF task actually display lower dynamical degeneracy. This reversed trend confirms that the changes in solution degeneracy arise from network size itself, not from residual variation in feature learning strength.
J.1 N-BFF
@@
J.3 Sine Wave Generation
@@
Appendix K Regularization’s effect on degeneracy for all tasks
@@
Appendix L Test feature learning effect on degeneracy in standard parameterization
While $\mu P$ lets us systematically vary feature-learning strength to study its impact on solution degeneracy, we confirm that the same qualitative pattern appears in standard discrete-time RNNs: stronger feature learning lowers dynamical degeneracy and raises weight degeneracy (Figure 26).
To manipulate feature-learning strength in these ordinary RNNs we applied the $\gamma$-trick—scaling the network’s outputs by $\gamma$—and multiplied the learning rate by the same factor. With width fixed, these two operations replicate the effective changes induced by $\mu P$. Figure 27 shows that this combination reliably tunes feature-learning strength. Besides weight-change norm and kernel alignment, we also report representation alignment (RA), giving a more fine-grained view of how much the learned features deviate from their initialization [53]. Representation alignment is the directional change of the network’s represenational dissimilarity matrix before and after training, and is defined by
@@
Appendix M Test network size effect on degeneracy in standard parameterization
@@
Appendix N Disclosure of compute resources

### v3

@@ 5 Acknowledgments
We acknowledge funding from NIH (RF1DA056403, U01NS136507), James S. McDonnell Foundation (220020466), Simons Foundation (Pilot Extension-00003332-02, McKnight Endowment Fund, CIFAR Azrieli Global Scholar Program, NSF (2046583), Harvard Medical School Neurobiology Lefler Small Grant Award, Harvard Medical School Dean’s Innovation Award, Alice and Joseph Brooks Fund Postdoctoral Fellowship, and Kempner Graduate Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University.
Appendix
@@ B.4 Path Integration
Appendix C Task performance of trained networks
In all experiments, we train networks until them reach a near‑asymptotic, task‑specific mean-squred error (MSE) threshold (0.001 for N-BFF, 0.01 for Delayed Discrimination, and 0.05 for Sine‑Wave Generation and Path Integration), after which we allow a patience period of 3 epochs and stop training to measure degeneracy. This early‑stopping criterion ensures that networks trained on the same task/condition achieve comparable final losses before any degeneracy analysis.
To quantify the residual variation, we report the coefficient of variation (CV) of the final training loss across seeds for each condition, expressed as % of the mean. Header labels match the x‑axis levels used in the main‑text figures. Final losses cluster tightly near small values of the loss threshold, so even a double-digit CV translates to very small absolute variation. For example, a 10% CV at an MSE of 0.001 implies an s.d. of $10^{-4}$; at 0.01 it’s $10^{-3}$. Additionally, the networks converged well on a global scale. Across our experiments, the mean MSE after training is under 2% of the mean MSE at initialization, indicating that training has converged well. Individual values: 0.059% (N-BFF), 1.6% (Delayed Discrimination), 0.32% (Sine-Wave Generation), 0.94% (Path Integration). CV can look large when the mean is tiny (the denominator is small). For example, a 16% CV on Sine-Wave Generation task corresponds to 0̃.05% of the initialization loss, which is consistent with minor differences due to the stochastic gradients rather than under-training.
These variability values are also not monotonic in any factor and sometimes move opposite to the degeneracy trends, arguing against a loss‑dispersion confound to solution degeneracy.
Table 2: Coefficient of variation (CV) of the final training loss across 50 networks for each task complexity level.
Task Complexity Level 1 Level 2 Level 3 Level 4
N-BFF 6.30% 4.60% 9.30% 3.50%
Delayed Discrim. 15.90% 8.40% 9.50%
Sine Wave Gen. 9.94% 9.20% 8.70%
Path Integr. 9.16% 2.85%
Table 3: Coefficient of variation (CV) of the final training loss across 50 networks for each feature learning strength ($\gamma$)
Feature Learning Strength $\boldsymbol{\gamma_{1}}$ $\boldsymbol{\gamma_{2}}$ $\boldsymbol{\gamma_{3}}$ $\boldsymbol{\gamma_{4}}$
N-BFF 9.70% 9.10% 13.40% 11.70%
Delayed Discrim. 8.70% 12.60% 11.70% 12.30%
Sine Wave Gen. 3.50% 3.90% 10.90% 11.70%
Path Integr. 5.40% 5.20% 6.20%
Table 4: Coefficient of variation (CV) of the final training loss across 50 networks for each network width.
Network Width 64 units 128 units 256 units
N-BFF 3.80% 4.20% 3.50%
Delayed Discrim. 3.30% 3.00% 3.20%
Sine Wave Gen. 17.80% 16.60% 16.40%
Path Integr. 5.10% 5.40% 5.90%
Table 5: Coefficient of variation (CV) of the final training loss across 50 networks for each L1 regularization strength.
L1 Regularization $\boldsymbol{\lambda_{1}}$ $\boldsymbol{\lambda_{2}}$ $\boldsymbol{\lambda_{3}}$ $\boldsymbol{\lambda_{4}}$
N-BFF 2.10% 6.90% 1.10%
Delayed Discrim. 15.90% 14.50% 16.70% 14.90%
Sine Wave Gen. 10.40% 11.10% 11.10%
Path Integr. 9.00% 7.10% 3.00%
Table 6: Coefficient of variation (CV) of the final training loss across 50 networks for each rank regularization strength.
Rank Regularization $\boldsymbol{\lambda_{1}}$ $\boldsymbol{\lambda_{2}}$ $\boldsymbol{\lambda_{3}}$ $\boldsymbol{\lambda_{4}}$
N-BFF 2.10% 7.20% 4.30%
Delayed Discrim. 15.90% 16.90% 13.60% 12.10%
Sine Wave Gen. 13.90% 14.30% 15.90%
Path Integr. 7.70% 7.90% 6.70%
Appendix D Memory demand of each task
@@
and train a two‐layer MLP to predict the subsequent target $y_{t+1}$. We then evaluate the held‐out mean‐squared error $\mathrm{MSE}(h)$, averaged over multiple random initializations. We identify the smallest history length $h^{*}$ at which the error curve plateaus or has a minimum, and take $h^{*}$ as the task’s intrinsic memory demand.
@@
Figure 9: Memory demand of each task. The held‐out mean‐squared error $\mathrm{MSE}(h)$ of a two‐layer MLP predictor is plotted against history length $h$. The intrinsic memory demand $h^{*}$, defined by the plateau or minimum of each curve, is 1 for the N‐Bits Flip-Flop and Path Integration tasks, 2 for Sine Wave Generation, and 25 for Delayed Discrimination—matching the inter‐stimulus delay interval in that task.
Appendix E Solution degeneracy in chaotic RNNs
Our original task suite comprises neuroscience‑motivated tasks that produce stable-attractors: fixed-point (N-Bit Flip Flop, Delayed Discrimination), limit cycle (Sine Wave Generation), and attractor manifold (Path Integration). To further demonstrate that the observed effect of the four factors on degeneracy extend to RNNs with chaotic activity, here we add a chaotic attractor task and verified that the effects of all four factors on dynamical and weight degeneracy are consistent with Table 1.
Lorenz 96 Attractor Dataset
We simulated trajectories from the Lorenz 96 dynamical system [lorenz1996predictability], defined by
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,\dots,N,$
with cyclic boundary conditions $x_{-1}=x_{N-1},\ x_{0}=x_{N},\ x_{N+1}=x_{1}$. The external forcing parameter was set to $F=8.0$, a standard choice that induces chaotic dynamics.
To generate the dataset, we numerically integrated the system for $N=16,24,$ and $32$ dimensions. Each simulation used a time step of $\Delta t=0.01$ and produced $15000$ time points after discarding an initial transient of $1000$ steps to remove non-stationary behavior. Initial conditions were sampled as small random perturbations around the fixed point $x_{i}=F$:
$x_{i}(0)=F+0.1\,\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,1).$
For each condition, we trained 50 RNNs on next-step prediction until the networks achieve a near-asymptotic MSE loss at $0.0005$. After training, the average Lyapunov exponent of RNNs trained on the Lorenz 96 attractor with 16 dimensions is $12.58\pm 0.74$, indicating chaotic neural dynamics.
Figure 10: RNN recurrent activities and Lorenz 96 attractor ($N=16$) trajectories projected onto their respective top 3 principle components.
Figure 11: Varying the four factors on Lorenz 96 next-step prediction task changes solution degeneracy across the dynamical and weight level in a way that is consistent with Table 1.
Appendix F Additional details on the degeneracy metrics
F.1 Dynamical Degeneracy
Briefly, DSA proceeds as follows: Given two RNNs with hidden states $\mathbf{h}_{1}(t)\in\mathbb{R}^{n}$ and $\mathbf{h}_{2}(t)\in\mathbb{R}^{n}$, we first generate a delay-embedded matrix, $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ of the hidden states in their original state space. Next, for each delay-embedded matrix, we use Dynamic Mode Decomposition (DMD) [schmid2022dynamic] to extract linear forward operators $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ of the two systems’ dynamics. Finally, a Procrustes distance between the two matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ is used to quantify the dissimilarity between the two dynamical systems and provide an overall DSA score, defined as:
@@ F.1 Dynamical Degeneracy
F.2 Weight degeneracy
We computed the pairwise distance between the recurrent matrices from different networks using Two-sided Permutation with One Transformation [schonemann1966generalized, ding2008nonnegative] function from the Procrustes Python package [meng2022procrustes].
F.3 Establishing a null distribution for dynamical and weight degeneracy
The DSA scores that we used to define the dynamical degeneracy are inherently context-dependent. Specifically, the absolute scale of DSA distances can vary with hyperparameters, particularly the delay embedding dimension and the rank used in DSA, because the underlying Procrustes analysis between two dynamics matrices relies on the Frobenius norm, which in turn depends on the dimension of the dynamic operator being compared. Following the procedure described in the original DSA paper, we fixed these hyperparameters across all groups within each task to ensure fair comparison.
To further validate the interpretation of DSA values, we computed null distributions of the DSA scores, i.e. the distribution of DSA scores when sampled neural activities come from identical networks. For each of the 50 networks analyzed in Figure 3B, we randomly split the sampled neural trajectories from the same network into two subsets and computed DSA distances between them. This procedure yields a distribution of DSA scores expected from identical dynamical systems, which serves as a reference noise floor. The 95% confidence intervals (CIs) for these null distributions are reported below (header labels such as "Level 1” correspond to the task-complexity levels shown in the main-text figures). These CIs are, on average, an order of magnitude smaller than the computed dynamical degeneracy, indicating that the observed differences between networks trained from different initializations are statistically significant.
Table 7: Establishing a null distribution for dynamical degeneracy: 95% confidence intervals of null DSA scores computed by comparing trajectories from the same network. CIs are on average an order of magnitude smaller than across-network distances.
Task Complexity Level 1 Level 2 Level 3 Level 4
N-BFF [0.011, 0.013] [0.009, 0.016] [0.008, 0.013] [0.006, 0.009]
Delayed Discrimination [0.039, 0.064] [0.014, 0.076] [0.025, 0.032]
Sine Wave Generation [0.057, 0.102] [0.054, 0.081] [0.048, 0.073]
Path Integration [0.023, 0.037] [0.010, 0.018]
For the PIF distance we used to define weight degeneracy, we similarly established a noise floor by randomly permuting each trained network’s recurrent weight matrix and computing the distance between the permuted and original matrices. The PIF metric reliably recovers a PIF distance of 0 under this null setting, confirming its robustness to noise and the meaningfulness of the reported cross-network PIF differences.
Appendix G Representational degeneracy
@@
$d_{\mathrm{repr}}(A_{x},A_{y})\;=\;1\;-\;\mathrm{SVCCA}\bigl(A_{x},A_{y}\bigr).$
@@
Appendix H Task complexity effect on degeneracy in Gated RNNs
To examine whether the observed trends in dynamical and weight degeneracy generalize beyond vanilla RNNs, we conducted additional experiments using gated recurrent units (GRUs). Note that prior work suggests that architectural choices influence the geometry but not the topology of neural dynamics, which is primarily shaped by task structure [gutenkunst2007universally]. Meanwhile, the Dynamical Similarity Analysis (DSA) metric we employ to quantify dynamical degeneracy is designed to precisely capture the topological organization of neural dynamics while remaining invariant to geometric transformations [ostrow_beyond_2023].
As a preliminary test, we trained GRUs on the Sine Wave Generation task while systematically varying task complexity by changing the number of input–output channels. Consistent with our findings in vanilla RNNs, increasing task complexity led to a decrease in dynamical degeneracy and a rise in weight degeneracy.
Figure 13: Increasing task complexity in the Sine Wave Generation task produces the same effect on dynamical and weight degeneracy in both vanilla RNNs and GRUs.
Appendix I Dense sweep on feature learning, network width, and regularization strength
it is important to know whether the degeneracy trends generalize to intermediate values and beyond the ranges reported in the main paper. To test this, we used 3-BFF as an example and ran a dense sweep both interpolating within and extrapolating beyond the ranges shown in Figs. 3 and 6–8. We demonstrate that cross dynamical and weight levels, the degeneracy trends remain consistent and interpolate smoothly across these intermediate values.
Figure 14: Feature learning, network width, and regularization strength’s effect on degeneracy over a denser sweep of conditions on the 3-Bits Flip Flop task.
Appendix J Detailed characterization of OOD generalization performance
@@ Appendix I Dense sweep on feature learning, network width, and regularization strength
J.1 Changing task complexity
@@
J.2 Changing feature learning strength
@@
J.3 Changing network size
@@
J.4 Changing regularization strength
J.4.1 Low-rank regularization
@@
Appendix K A short introduction to Maximal Update Parameterization ($\mu P$)
Under the NTK parametrization, as the network width goes to infinity, the network operates in the lazy regime, where its functional evolution is well‑approximated by a first‑order Taylor expansion around the initial parameters [jacot2018neural, lee2019wide, chizat2019lazy, woodworth2020kernel]. In this limit feature learning is suppressed and training dynamics are governed by the fixed Neural Tangent Kernel (NTK).
[…]
@@
Appendix L Theoretical relationship between parameterizations
[…]
@@
[…]
Miller and Fumarola [miller2012mathematical] demonstrated that two classes of continuous-time firing-rate models which differ in their placement of the non-linearity are mathematically equivalent under a change of variables:
@@
[…]
@@
In summary, while a theoretical equivalence exists, it is contingent on consistent scaling across all components of the model. In this manuscript, we use the standard discrete-time RNNs due to its practical relevance for task-driven modeling community, while switching to $\mu P$ to isolate the effect of feature learning and network size. Additionally, we confirm that the feature learning and network size effects on degeneracy hold qualitatively the same in standard discrete-time RNNs, unless where altering network width induces unstable and lazier learning in larger networks (Figure P and Q).
Appendix M Verifying larger $\gamma$ reliably induces stronger feature learning in $\mu P$
In $\mu P$ parameterization, the parameter $\gamma$ interpolates between lazy training and rich, feature‐learning dynamics, without itself being the absolute magnitude of feature learning. Here, we assess feature‐learning strength in RNNs under varying $\gamma$ using two complementary metrics:
@@
M.3 Sine Wave Generation
@@
Appendix N Verifying $\mu P$ reliably controls for feature learning across network width
Here, we only use Kernel Alignment to assess the feature learning strength in the networks since the unnormalized weight‐change norm $\left\|\mathbf{W}_{T}-\mathbf{W}_{0}\right\|_{F}$ scales directly with matrix size (therefore network size) and there exists no obvious way to normalize across different dimensions. In our earlier analysis where we compared weight‐change norms at varying $\gamma$, network size remained fixed, so those Frobenius‐norm measures were directly comparable. We found that, for all tasks except Delayed Discrimination, the change in mean KA across different network sizes remains extremely small (less than 0.1), which demonstrates that $\mu P$ parameterization with the same $\gamma$ has effectively controlled for feature learning strength across network sizes. On Delayed Discrimination, the networks undergo slightly lazier learning for larger network sizes. Nevertheless, we still include Delayed Discrimination in our analyses of solution degeneracy to ensure our conclusions remain robust even when $\mu P$ can’t perfectly equalize feature‐learning strength across widths. As shown in the main paper, lazier learning regime generally increases dynamical degeneracy; yet, larger networks which exhibit lazier learning in the N-BFF task actually display lower dynamical degeneracy. This reversed trend confirms that the changes in solution degeneracy arise from network size itself, not from residual variation in feature learning strength.
N.1 N-BFF
@@
N.3 Sine Wave Generation
@@
Appendix O Regularization’s effect on degeneracy for all tasks
@@
Appendix P Test feature learning effect on degeneracy in standard parameterization
While $\mu P$ lets us systematically vary feature‑learning strength to study its impact on solution degeneracy, we confirm that the same qualitative pattern appears in standard discrete‑time RNNs: stronger feature learning lowers dynamical degeneracy and raises weight degeneracy (Figure 30).
To manipulate feature‑learning strength in these ordinary RNNs we applied the $\gamma$‑trick—scaling the network’s outputs by $\gamma$—and multiplied the learning rate by the same factor. With width fixed, these two operations replicate the effective changes induced by $\mu P$. Figure 31 shows that this combination reliably tunes feature‑learning strength. Besides weight‑change norm and kernel alignment, we also report representation alignment (RA), giving a more fine‑grained view of how much the learned features deviate from their initialization [liu2023connectivity]. Representation alignment is the directional change of the network’s representational dissimilarity matrix before and after training, and is defined by
@@
Appendix Q Test network size effect on degeneracy in standard parameterization
@@
Appendix R Disclosure of compute resources